OpenStudy (anonymous):

Given \(\vec{g} = -3\vec{i} + 6\vec{j} - 12\vec{k}\) and \(\vec{h} = 5\vec{i} - 10\vec{j} + 20\vec{k}\), how could I prove that \(\vec{g}\) and \(\vec{h}\) are parallel? Please explain the reasoning.

OpenStudy (anonymous):

take their vector product, or prove they are multiple of each other

OpenStudy (anonymous):

Well does a vector product of 0 allow me to conclude that they are parallel?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

g x h=|g||h|sin(theta) where theta is angle between them. It will be = 0 only if angle is 0º or 180º

OpenStudy (anonymous):

Ok, just clarifying, is it because using the formal definition of the magnitude of a vector product, if \(\theta\) = 0º or 180º, it would cause the entire thing to become 0, and therefore parallel since an angle of 0º or 180º would be the same line or parallel?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh, never mind, thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

more easy: -5/3(-3i+6j-12h)=5i-10j+20k

OpenStudy (anonymous):

so -5/3g=h

OpenStudy (anonymous):

Ok. that is the reason the book gave me, so I'm curious as to why that works...I don't see where -5/3 comes from.

OpenStudy (anonymous):

to make -3 equal to 5

OpenStudy (anonymous):

Oh. so you're multiplying \(\vec{g}\) by -5/3 to make it equal \(\vec{h}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, once again, thank you :)

OpenStudy (anonymous):

yw, :)